OpenStudy (anonymous):

open parentheses 6 plus 3 square root of 2 close parentheses open parentheses 5 minus square root of 2 close parentheses

OpenStudy (anonymous):

u meant \[(6+3\sqrt{2}) (5-\sqrt{2}) \] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

"distribute" the 6, then "distribute" the 3 root 2 and what do you get?

OpenStudy (anonymous):

is it 30 + 6 sqrt 2?

OpenStudy (anonymous):

expand the bracket u'll get \[30- 6\sqrt{2} + 15\sqrt{2 - 6}\] which will give you z answer \[ 24 -9\sqrt{2}\] or \[8- 3\sqrt{2}\] when divided by 3

OpenStudy (anonymous):

that was half of it, but you missed a sign

OpenStudy (anonymous):

24 + 9 sqrt 2?

OpenStudy (anonymous):

good job... can you simplify that?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

dont 24 and 9 have a common factor?

OpenStudy (anonymous):

you got the right answer, but they tend to like to see 3 ( 8 + 3 root 2) for some reason.... but either is ok... good job